OpenStudy (anonymous):

Find the radius of the circle x^2+y^2=5

OpenStudy (agent0smith):

\[x^2 + y^2 = 5\]standard form for a circle is \[x^2 +y^2 = r^2\]where r is the radius. In your case.. \[r^2 = 5\]

OpenStudy (anonymous):

Ohhh..ok..thank you :)

Directrix (directrix):

Center of the given circle is at the origin: (0,0) Standard Form: (x - 0) ^ 2 + (y - 0) ^ 2 = (?)^2